# Generalized Friendship Paradoxes in Network Science

### v2

@@ Abstract
Generalized friendship paradoxes occur when, on average, our friends have more of some attribute than us. These paradoxes are relevant to many aspects of human interaction, notably in social science and epidemiology. Here, we derive new theoretical results concerning the inevitability of a paradox arising, using a linear algebra perspective. Following the seminal 1991 work of Scott L. Feld, we consider two distinct ways to measure and compare averages, which may be regarded as global and local. For global averaging, we show that a generalized friendship paradox holds for a large family of walk-based centralities, including Katz centrality and total subgraph communicability, and also for nonbacktracking eigenvector centrality. Defining loneliness as the reciprocal of the number of friends, we show that for this attribute the generalized friendship paradox always holds in reverse. In this sense, we are always more lonely, on average, than our friends. We also derive a global averaging paradox result for the case where the arithmetic mean is replaced by the geometric mean. The nonbacktracking eigenvector centrality paradox is also established for the case of local averaging.
@@ 1 Motivation
The original work of Feld [17] set up two distinct types of averaging process. Following [21] we will distinguish between them by using the terms global and local. (We note that the terms “ego” and “alter” are used in [36].) Among the empirical tests for paradox effects across various application domains, some emphasize local averaging [2, 12, 6, 9, 34, 40, 53, 54] and others global averaging [25, 28, 41]. In this work we contribute to the understanding of both types of generalized friendship paradox by highlighting new cases where a paradox is guaranteed to hold for any network. We thereby add to the theoretical inevitability results on global averaging in [17, 27] and on local averaging in [9, 14, 26]. We make several contributions:
@@
  • Theorem 7.1 shows that a global geometric mean paradox holds for degree.
@@ 2.1 Global Averages
We will refer to (3) as a generalized global friendship inequality; if this holds then, on average, our neighbours have more of the attribute $\mathbf{x}$ than we do. We emphasize that in the general case where the attribute vector $\mathbf{x}$ is independent of the network structure, there is no way of knowing a priori whether the inequality (3) will hold. However, there is a simple and intuitive characterisation: Eom and Jo [14] showed that the inequality is equivalent to the statement that $\mathbf{x}$ is nonnegatively correlated with the degree vector.
@@ 2.2 Local Averages
$\mathbf{x}=A^{\ell}\mathbf{1},$ (7)
where $\ell>1$ is an integer, and for Katz centrality [33], defined as
@@
The generalized local friendship inequality (5) holds for eigenvector, walk and Katz centrality, with equality if and only if the network is regular.
We also note that the analysis in [26] establishes equality if and only if the network is regular in the cases of eigenvector and Katz centrality.
In the following two sections we demonstrate that further global and local paradoxes may be established.
@@ 3.1 Generalized Global Friendship Paradox for Katz-style Centralities
The generalized global friendship inequality (3) holds for $\mathbf{x}=f(A)\mathbf{1}$, where $f$ is an increasing function defined over the spectrum of $A$. (In particular, it holds when $f$ has a convergent power series with positive coefficients.) Furthermore, if $f$ is strictly increasing then equality holds if and only if the associated graph is regular.
@@
Because $f$ is increasing we have $(\lambda_{i}-\lambda_{j})(f(\lambda_{i})-f(\lambda_{j}))q\geq 0$ for any $i$ and $j$. Summing, we see that the expression on the right hand side of (13) is greater than or equal to zero, and hence (12) holds.
@@
Theorem 3.1 immediately covers walk centrality (7), Katz centrality (8) and total subgraph communicability [3] (also known as exponential centrality), which is defined by
@@ 3.2 Generalized Global Friendship Paradox for Nonbacktracking Eigenvector Centrality
For some networks, the standard eigenvector centrality measure has been observed to exhibit localization—attributing large values to a small subset of very central nodes, and failing to distinguish adequately between the remainder [1, 35, 38, 44]. An alternative nonbacktracking version of eigenvector centrality was proposed in [38].
@@
$\left[\begin{array}[]{cc}A&I-D\\ I&0\end{array}\right]\left[\begin{array}[]{c}\mathbf{v}_{a}\\ \mathbf{v}_{b}\end{array}\right]=\lambda\left[\begin{array}[]{c}\mathbf{v}_{a}\\ \mathbf{v}_{b}\end{array}\right].$ (15)
Here, it is known that a dominant eigenvalue $\lambda$ of this $2n$ by $2n$ matrix exists, that is real-valued and positive, and in the corresponding eigenvector both $\mathbf{v}_{a}$ and $\mathbf{v}_{b}$ can be chosen to have nonnegative entries [24, 38]. As suggested in [38], we may then take $\mathbf{v}_{a}$ to be the nonbacktracking eigenvector centrality measure. We will normalize $\mathbf{v}_{a}$ so that $\mathbf{1}^{T}\mathbf{v}_{a}=1$.
The eigenvalue problem (15) is intimately connected with the nonbacktracking, or Hashimoto, matrix [38, 43, 48] and also with nonbacktracking Katz-style centrality [24, Theorem 10.2].
@@
First we note that $\lambda\geq 1$, since
$\left[\begin{array}[]{cc}A&I-D\\ I&0\end{array}\right]\left[\begin{array}[]{c}\mathbf{1}\\ \mathbf{1}\end{array}\right]=\left[\begin{array}[]{c}\mathbf{1}\\ \mathbf{1}\end{array}\right].$
If $\lambda=1$ then $\mathbf{v}_{a}=\mathbf{1}$ and the inequality holds with equality.
Hence, we now assume that $\lambda>1$.
@@
We now make use of the connection between the eigenvalue problem (15), the nonbacktracking matrix and the radius of convergence, $R$, of the Ihara zeta function of the graph. It is known [29, Thm 5.6.15 applied to the matrix $zA$] that $\lambda$ is equal to the largest eigenvalue of the nonbacktracking matrix, which itself is equal to the reciprocal of $R$; so $\lambda=1/R$. Terras [50] conjectured that
@@
It is shown in [22, Corollary 3.2.1] that increasing the size of the graph by taking one node and joining it to a tree does not change $\lambda$. Hence, removing a single node of degree one, along with the corresponding edge, does not change $\lambda$. See also [43, Theorem 1.2] for a more direct argument that also applies to directed graphs. We may therefore iteratively remove all nodes of degree one until we have a graph where all nodes have degree at least two, whence the analysis above applies if the remaining graph is not a cycle.
If the graph is a cycle, then from [51, Proposition 3.6] we have $\lambda=1$, and hence, as argued at the start of the proof, equality arises. ∎
@@ 6 Loneliness Paradoxes
We use similar arguments to those in [9, Section 2]. Taking $x_{i}=1/d_{i}$, in the left hand side of (21) we have
$\mathbf{1}^{T}(D^{-1}A\mathbf{x}-\mathbf{x})=\sum_{i,j}\left(\frac{a_{i,j}}{d_{i}d_{j}}\right)-\left(\sum_{i}\frac{1}{d_{i}}\right)=\sum_{i,j}\left(\frac{a_{i,j}}{d_{i}d_{j}}-\frac{a_{i,j}}{d_{i}^{2}}\right)=\sum_{i,j}\left(a_{i,j}\left(\frac{1}{d_{i}d_{j}}-\frac{1}{d_{i}^{2}}\right)\right).$
@@
and the result follows.
7 A Global Geometric Mean Friendship Paradox
In this section we will make use of the following lemma.
@@ 7 A Global Geometric Mean Friendship Paradox
$\sum_{i}\sum_{j}\left(d_{i}-d_{j}\right)\left(f(d_{i})-f(d_{j})\right)\geq 0,$
with equality if and only if the network is regular. Upon simplification and rearrangement, this inequality reduces to
@@
The aim of this section is to consider using the geometric mean in (2) rather than the arithmetic mean. As mentioned in the start of subsection 2.1, the friend-of-friend list includes each degree value $d_{i}$ exactly $d_{i}$ times, and the list involves $\mathbf{1}^{T}\mathbf{d}$ numbers. So the friend-of-friend geometric mean is

### v3

@@ Abstract
Generalized friendship paradoxes occur when, on average, our friends have more of some attribute than us. These paradoxes are relevant to many aspects of human interaction, notably in social science and epidemiology. Here, we derive new theoretical results concerning the inevitability of a paradox arising, using a linear algebra perspective. Following the seminal 1991 work of Scott L. Feld, we consider two distinct ways to measure and compare averages, which may be regarded as global and local. For global averaging, we show that a generalized friendship paradox holds for a large family of walk-based centralities, including Katz centrality and total subgraph communicability, and also for nonbacktracking eigenvector centrality. However, we also find counterexamples for centralities based on walks of even length. For local averaging we establish a paradox for nonbacktracking eigenvector centrality and we characterize the cases where the paradox holds with equality for the walk-based case. Defining loneliness as the reciprocal of the number of friends, we show that for this attribute the generalized and local friendship paradoxes always hold in reverse. In this sense, we are always more lonely, on average, than our friends. We also derive global and local averaging paradoxes for the case where the arithmetic mean is replaced by the geometric mean. As well as unifying and adding to the literature in this area, we highlight some open questions.
@@ 1 Motivation
The original work of Feld [24] set up two distinct types of averaging process. Following [26] we will distinguish between them by using the terms global and local. (We note that the terms “ego” and “alter” are used in [47].) Among the empirical tests for paradox effects across various application domains, some emphasize local averaging [4, 14, 7, 10, 45, 51, 69, 71] and others global averaging [33, 38, 52]. In this work we contribute to the understanding of both types of generalized friendship paradox by highlighting new cases where a paradox is guaranteed to hold for any network. We thereby add to the theoretical inevitability results on global averaging in [24, 37] and on local averaging in [10, 18, 36]. Our main contributions are:
@@
  • Theorems 7.1 and 7.2 show that global and local geometric mean paradoxes hold for degree.
  • Theorem 8.1 characterizes when equality holds in the local averaging paradox for walk centrality.
  • Counterexamples in subsection 8.2 show that a paradox does not hold with global averaging for even length walk centrality.
Section 2 sets up the key inequalities and describes known results. In section 37 we derive new paradoxes, and in section 8 we deal with some extra technicalities. The results of computational experiments on real and synthetic networks are presented in section 9. In section 10 we summarize the current state of the art and highlight some open questions.
@@ 2.1 Global Averages
We will refer to (3) as a generalized global friendship inequality; if this holds then, on average, our neighbours have more of the attribute $\mathbf{x}$ than we do. We emphasize that in the general case where the attribute vector $\mathbf{x}$ is independent of the network structure, there is no way of knowing a priori whether the inequality (3) will hold. However, there is a simple and intuitive characterisation: Eom and Jo [18] showed that the inequality is equivalent to the statement that $\mathbf{x}$ is nonnegatively correlated with the degree vector. (We say that the vectors $\mathbf{u}$ and $\mathbf{v}$ in $\mathbb{R}^{n}$ are nonnegatively correlated if $\sum_{i=1}^{n}\sum_{j=1}^{n}(u_{i}-u_{j})(v_{i}-v_{j})\geq 0$, and positively correlated if the inequality holds strictly.)
@@ 2.2 Local Averages
$\mathbf{x}=A^{\ell}\mathbf{1},\quad\text{~where~}\ell>1\text{~is~an~integer},$ (7)
and for Katz centrality [43], defined as
@@
The generalized local friendship inequality (5) holds for eigenvector, walk and Katz centrality.
The authors in [36] also discuss conditions under which equality holds in (5). In the case of eigenvector and Katz centrality, they prove that equality holds if and only if the graph is regular. For walk centrality, they point out that equality always holds in the degenerate case where $\ell=0$ in (7), so $\mathbf{x}=\mathbf{1}$. However, equality is not discussed for $\ell\geq 1$. We return to this issue in section 8, where we show in Theorem 8.1 that for odd $\ell$ equality holds if and only if the graph is regular and for even $\ell$ equality holds if and only if the graph is either regular or biregular.
In the following sections we demonstrate that further global and local paradoxes may be established.
Remark 2.1.
We note that the inequalities (3) and (5) have two natural properties. First, if $\mathbf{x}$ satisfies an inequality then so does $\beta\mathbf{x}$ for any $\beta>0$. Second, if $\mathbf{x}$ and $\mathbf{y}$ satisfy an inequality, then so does $\mathbf{x}+\mathbf{y}$.
@@ 3.1 Generalized Global Friendship Paradox for Katz-style Centralities
The generalized global friendship inequality (3) holds for $\mathbf{x}=f(A)\mathbf{1}$, where $f$ is an increasing function defined over the spectrum of $A$. Furthermore, if $f$ is strictly increasing then equality holds if and only if the associated graph is regular.
@@
Because $f$ is increasing we have $(\lambda_{i}-\lambda_{j})(f(\lambda_{i})-f(\lambda_{j}))\geq 0$ for any $i$ and $j$. Summing, we see that the expression on the right hand side of (13) is greater than or equal to zero, and hence (12) holds.
@@
Theorem 3.1 immediately covers walk centrality (7), Katz centrality (8) and total subgraph communicability [5], also known as exponential centrality, which is defined by
@@
Note that the generalized global friendship inequality does not hold in general for the centrality measure given by walk centrality when $\ell$ is even. We discuss this issue in section 8.
@@ 3.2 Generalized Global Friendship Paradox for Nonbacktracking Eigenvector Centrality
For some networks, the standard eigenvector centrality measure has been observed to exhibit localization—attributing large values to a small subset of very central nodes, and failing to distinguish adequately between the remainder [2, 46, 49, 58]. To remedy this, an alternative nonbacktracking version of eigenvector centrality was proposed in [49].
@@
$\left[\begin{array}[]{cc}A&I-D\\ I&0\end{array}\right]\left[\begin{array}[]{c}\mathbf{v}_{a}\\ \mathbf{v}_{b}\end{array}\right]=\lambda\left[\begin{array}[]{c}\mathbf{v}_{a}\\ \mathbf{v}_{b}\end{array}\right],$ (15)
which is intimately connected with the nonbacktracking, or Hashimoto, matrix [49, 56, 63] and also with nonbacktracking Katz-style centrality [32, Theorem 10.2]. Generically, the $2n$ by $2n$ matrix in (15) has a spectral radius strictly greater than one, with a dominant real-valued eigenvalue $\lambda>1$ for which $\mathbf{v}_{a}$ and $\mathbf{v}_{b}$ in the corresponding eigenvector can be chosen to have nonnegative entries. Following the authors in [49], we then use $\mathbf{v}_{a}$ to define the nonbacktracking eigenvector centrality measure, normalized so that $\mathbf{1}^{T}\mathbf{v}_{a}=1$. However, for certain graphs the spectral radius is equal to one; this is a case where the dominant eigenvalue may not be not unique. It is immediate in (15) that $\lambda=1$ is an eigenvalue with corresponding eigenvector given by $\mathbf{v}_{a}=\mathbf{v}_{b}=\mathbf{1}$. Hence when $\lambda=1$ is a dominant eigenvalue we take the natural choice of $\mathbf{x}=\mathbf{1}/n$ to define the centrality measure. In this scenario, where the centrality measure is uniform, both versions of the generalized friendship inequality hold with equality. So it is of interest to characterize when $\lambda=1$ is a dominant eigenvalue. This motivates the material in subsection 8.6. For completeness, we also provide arguments in subsection 8.7 which confirm that the centrality measure is always well defined.
@@
Let $\lambda\geq 1$ denote the largest real-valued eigenvalue of the system (15). As mentioned above, if $\lambda=1$ then $\mathbf{v}_{a}=\mathbf{1}/n$ and the inequality holds with equality.
Hence we may assume that there is a unique real dominant eigenvalue $\lambda>1$.
@@
We now make use of the connection between the eigenvalue problem (15), the Hashimoto, or nonbacktracking, matrix $B$ (which we define in subsection 8.6), and the radius of convergence, $R$, of the Ihara zeta function of the graph. The Ihara zeta function is defined as
$\zeta(u)^{-1}=\mathrm{det}(I-uB).$
It is known that $\lambda$ is equal to the largest eigenvalue of the nonbacktracking matrix, which itself is equal to the reciprocal of $R$ (see [39, Thm 5.6.15] applied to the matrix $zA$); so $\lambda=1/R$. Terras [65] conjectured that
@@
It is shown in [30, Corollary 3.2.1] that increasing the size of the graph by taking one node and joining it to a tree does not change $\lambda$. Hence, removing a single node of degree one, along with the corresponding edge, does not change $\lambda$. See also [56, Theorem 1.2] for a more direct argument that also applies to directed graphs. We may therefore iteratively remove all nodes of degree one until we have a graph where all nodes have degree at least two, whence the analysis above applies if the remaining graph is not a cycle.
If the graph is a cycle, then from [66, Proposition 3.6] $\lambda=1$ and we return to the equality case. ∎
@@ 6 Loneliness Paradoxes
We use similar arguments to those in [10, Section 2]. Taking $x_{i}=1/d_{i}$, in the left hand side of (21) we have
$\mathbf{1}^{T}(D^{-1}A\mathbf{x}-\mathbf{x})=\sum_{i,j}\left(\frac{a_{ij}}{d_{i}d_{j}}\right)-\left(\sum_{i}\frac{1}{d_{i}}\right)=\sum_{i,j}\left(\frac{a_{ij}}{d_{i}d_{j}}-\frac{a_{ij}}{d_{i}^{2}}\right)=\sum_{i,j}\left(a_{ij}\left(\frac{1}{d_{i}d_{j}}-\frac{1}{d_{i}^{2}}\right)\right).$
@@
giving the required inequality. We see that equality requires $d_{i}=d_{j}$ whenever nodes $i$ and $j$ share an edge. Since the graph is connected, this requires the graph to be regular. ∎
7 Geometric Mean Friendship Paradoxes
The aim of this section is to consider using the geometric mean in (2) and (4) rather than the arithmetic mean.
7.1 Global Geometric Mean Paradox
To study the global case, we will make use of the following lemma.
@@ 7.1 Global Geometric Mean Paradox
$\sum_{i,j}\left(d_{i}-d_{j}\right)\left(f(d_{i})-f(d_{j})\right)\geq 0,$
with equality if and only if the network is regular. Hence, $f(\mathbf{d})$ is nonnegatively correlated with $\mathbf{d}$, and the correlation is positive unless the graph is regular.
Upon simplification and rearrangement, this inequality reduces to
@@
As mentioned in the start of subsection 2.1, the friend-of-friend list includes each degree value $d_{i}$ exactly $d_{i}$ times, and the list involves $\mathbf{1}^{T}\mathbf{d}$ numbers. So the friend-of-friend geometric mean is